OpenStudy (anonymous):

Find the derivative of f(x) = -9/x at x = -8.

OpenStudy (p0sitr0n):

f'(x) = 9/x^2 since f'(u(x))=u' * x' f'(8)=9/64

OpenStudy (anonymous):

that makes sense, thanks!

OpenStudy (p0sitr0n):

no prob

OpenStudy (anonymous):

can you help me on one more question?